OpenStudy (anonymous):

Check: What are the directrices of the hyperbola given by the equation x^2/25-y^2/81=1? My Answer: x=±(25√106)/106 Option 2: y=±(25√106)/106 Option 3: x=±(81√106)/106 Option 4: y=±(81√106)/106

OpenStudy (anonymous):

hmm.... are you sure those are the correct choices?

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

oh... sorry... directrices... i was thinking asymptotes..

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

since the directrix is \(\large x=\frac{a^2}{c} \) did you find the value for c?

OpenStudy (anonymous):

106 right?

OpenStudy (anonymous):

well \(\large c^2=106\rightarrow c=\sqrt{106} \)

OpenStudy (anonymous):

yes... your answer looks right....:)

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

yw...:)